# Learning Versatile Optimizers on a Compute Diet

### v1

@@ 8 Scope and limitations
Our work focuses on a recipe to train learned optimizers that meta-generalize well from a limited training set. However, this work is leaving out a lot of “free” performance gains which could be obtained by trivially modifying or expanding the meta-training set, using better/faster architectures like state-space models instead of LSTMs. We intentionally kept our architecture and meta-training setup simple without any bells and whistles to do a fair head-to-head comparison with prior work and focus solely on the improvements from our recipe. Moreover, although our work significantly improves generalization performance of fully learned optimizers and we firmly believe this work is the right step towards making learned optimizers more practical, it is not focused on delivering the most-performant optimizer for all kinds of large-scale workloads. Additional work needs to be done to make these learned optimizers ready for production use cases like analyzing the impact of meta-training data distribution on generalization properties which is out of scope for this work. Furthermore, given the amount of compute resources we had, we tried our best to include as diverse and large number of tasks as possible for evaluating all the learned optimizers and ablations with multiple seeds and hyper-parameters. This resulted in hundreds of evaluation runs for each optimizer ablation. However, there is a scope for larger scale study just focused on evaluation using our proposed metrics which we leave for future work.
@@ 9 Conclusion
10 Acknowledgements
We acknowledge support from Mila-Samsung Research Grant, FRQNT New Scholar [EB], the FRQNT Doctoral (B2X) scholarship [AM] and the Canada-CIFAR AI Chair program [GL]. We also acknowledge resources provided by Compute Canada, Calcul Québec, and Mila.
@@ A.2 Additional results
In order to be fair in comparison with all the RNN-based learned optimizer baselines which use, we scale the hidden size of VeLO from 512 to 64 and accordingly the MLP bank size from 256 to 32 as mentioned in §5. In Table 5, we show meta-generalization results with original VeLO’s original architecture (denoted by “velo”) meta-trained with the same setup as other baselines (§5). Morever, we also show results for the released pre-trained VeLO model Metz et al. (2022b) (denoted by “pretrained-velo” in Table 5) which is meta-trained using 4000 months on large collection of tasks. Note that for pre-trained VeLO, our evaluation task set is actually in-distribution (Metz et al., 2022b) but for VeLO, VeLO-S and also our proposed Celo optimizer, they are out-of-distribution. The meta-generalization performance of VeLO is slightly better with the bigger hidden size (512) than VeLO-S (IQM 0.96 vs 0.90) but still lags behind our proposed approach Celo (IQM 1.20). This result shows that returns from architecture and algorithmic improvements in learned optimizers can be much higher than simply scaling optimizer sizes.
@@ A.3 Implementation details
A.4 Meta-training tasks
@@ A.6 Evaluation plots
A.7 Evaluation plots with task augmentation

### v2

@@ 8 Limitations and future work
Our work focuses on a recipe to train learned optimizers that meta-generalize well from a limited training set. However, this work is leaving out a lot of “free” performance gains which could be obtained by trivially modifying or expanding the meta-training set, using better/faster architectures like state-space models instead of LSTMs. We intentionally kept our architecture and meta-training setup simple without any bells and whistles to do a fair head-to-head comparison with prior work and focus solely on the improvements from our recipe. Moreover, although our work significantly improves generalization performance of fully learned optimizers and we firmly believe this work is the right step towards making learned optimizers more practical, it is not focused on delivering the most-performant optimizer for all kinds of large-scale workloads. Additional work needs to be done to make these learned optimizers ready for production use cases like analyzing the impact of meta-training data distribution on generalization properties, which is out of scope for this work. Furthermore, given the amount of compute resources we had, we tried our best to include as diverse and large number of tasks as possible for evaluating all the learned optimizers and ablations with multiple seeds and hyperparameters. This resulted in hundreds of evaluation runs for each optimizer ablation. However, there is a scope for larger scale study just focused on evaluation using our proposed metrics which we leave for future work.
@@ 9 Conclusion
Acknowledgements
We acknowledge support from Mila-Samsung Research Grant, FRQNT New Scholar [EB], the FRQNT Doctoral (B2X) scholarship [AM] and the Canada-CIFAR AI Chair program [GL]. We also acknowledge resources provided by Compute Canada, Calcul Québec, and Mila. We thank NVIDIA for providing the GPUs that made this work possible. We are also grateful to the developers of open-source libraries such as JAX (Bradbury et al., 2018), NumPy (Harris et al., 2020), google/learned_optimization, and google-research/rliable, which were instrumental in this research.
@@ A.2 Additional results
In order to be fair in comparison with all the RNN-based learned optimizer baselines which use, we scale the hidden size of VeLO from 512 to 64 and accordingly the MLP bank size from 256 to 32 as mentioned in §5. In Table 5, we show meta-generalization results with original VeLO’s original architecture (denoted by “velo”) meta-trained with the same setup as other baselines (§5). Moreover, we also show results for the released pre-trained VeLO model Metz et al. (2022b) (denoted by “pretrained-velo” in Table 5) which is meta-trained using 4000 months on large collection of tasks. Note that for pre-trained VeLO, our evaluation task set is actually in-distribution (Metz et al., 2022b) but for VeLO, VeLO-S and also our proposed Celo optimizer, they are out-of-distribution. The meta-generalization performance of VeLO is slightly better with the bigger hidden size (512) than VeLO-S (IQM 0.96 vs 0.90) but still lags behind our proposed approach Celo (IQM 1.20). This result shows that returns from architecture and algorithmic improvements in learned optimizers can be much higher than simply scaling optimizer sizes.
@@
Task augmentation level ablations.
Task augmentation i.e. re-parametrization can be applied at different hierarchical levels of neural network during meta-training: (1) Global level, where a single augmentation parameter is sampled for the entire network; (2) Tensor level, where an augmentation parameter is sampled for each tensor or layer in the network; (3) Per-parameter level, where an augmentation parameter is sampled individually for each parameter in the network; and (4) Mixed strategy, where the augmentation parameter is sampled uniformly from a combination of the previous three approaches. In our experiments, for mixed augmentation strategy, we first randomly sample a strategy from the list ["none", "none", "global", "global", "tensor", "tensor", "parameter"] and then randomly select an augmentation range from the list [[0.001, 1000.0], [0.01, 100.0], [0.1, 10.0]]. Additionally, different augmentation strategies can used in the two stages for our approach presented in §3. We present results for all the ablations in Table 6. Overall, we find that global augmentation performs the best closely followed by tensor augmentation.
task aug level
[…]
tensor global 1.13 0.04
global tensor 1.13 0.04
mix mix 1.14 0.03
tensor tensor 1.18 0.04
celo global global 1.20 0.01
Table 6: Task augmentation level ablations. IQM and optimality gap (OG) metrics for final loss criteria on our 17 evaluation task set (§5) are reported. Task augmentation is applied at different levels during two-stage meta-training proposed in §3. "mix" refers to randomly sampling from a mixture of level list, see § A.2 for more details.
@@ A.3 Implementation details
A.4 Scaling study
In our main paper, learned optimizers are meta-trained with a 2K unroll length. To study scaling behavior of our proposed approach, we meta-train Celo and VeLO-S with unroll lengths of 1K, 2K, and 5K, then meta-test them on all tasks in our evaluation set using a 10K unroll length, assessing generalization to longer unrolls. We report the IQM metric with the final loss criterion and show scaling trends in the right Figure. Celo performance scales predictably with the meta-training budget and gets IQM 1.85 on 10K unroll length with 5K meta-train unroll length, hence demonstrating strong generalization performance with respect to unroll length. VeLO-S, on the other hand, gets IQM 0.34 with 5K meta-train length. These results clearly demonstrate that Celo scales significantly better than VeLO, achieving much better performance under a fixed meta-training budget, thereby highlighting its compute efficiency.
A.5 Runtime analysis
Celo updates the parameters of a given optimizee network using a per-parameter MLP. Since the number of FLOPs for a forward pass of a linear layer with weights $n\times m$ is $2nm$, the cost of a 2-layer MLP update rule in Celo with 30 input features, 2 hidden layers with 4 units and 2 outputs units is $2\times(30\times 4+4\times 4+4\times 2)=288$ FLOPs per parameter. For an optimizee linear layer with $n\times m$ parameters and batch size $B$, the total amount of FLOPs to perform a forward and backward pass is $6nmB$. Hence, the ratio of FLOPs of our learned optimizer’s update w.r.t. to forward+backward FLOPs is $288nm/6nmB=48/B$. This shows that the overhead of our optimizer decreases as the tasks get larger with more inputs flowing through the optimizee network with larger batch size $B$ which also makes sense intuitively because as the tasks become computationally expensive, the cost of forward and backward pass dominates optimizer update cost. In this analysis, we omitted the cost of LSTM scheduler but as noted by prior work (Metz et al., 2022b; a), the overhead of any tensor or global-level computation in the learned optimizer decreases as the number of parameters increase in the optimizee network since the cost of per-parameter updates dominate and tensor/global computation overhead asymptotically reduces to a fixed constant (Metz et al., 2022b). We compare wall-clock time of Celo with Adam and other learned optimizer baselines in Table 7. We benchmark runtime on a NVIDIA V100 GPU with 10 seeds per optimizer.
runtime per step (ms)
evaluation tasks adam velo-s velo celo
ImageMLP_FashionMnist_Relu128x128 0.20 1.69 1.76 1.31
ImageMLP_Cifar10_128x128x128_LayerNorm_Relu 0.31 3.41 3.83 2.69
ImageMLP_Cifar10_128x128x128_Tanh_bs128 0.24 2.74 3.05 2.20
Conv_Cifar10_32x64x64 1.12 2.80 3.02 2.44
Conv_Cifar10_32x64x64_batchnorm 1.29 3.68 4.10 3.22
Conv_Cifar10_32x64x64_layernorm 1.41 3.78 4.11 3.21
Conv_Cifar100_32x64x64 1.11 2.70 3.09 2.55
VIT_Cifar100_wideshallow 6.73 43.89 46.00 39.47
VIT_Cifar100_skinnydeep 4.52 29.09 32.69 23.70
ImageMLPAE_Cifar10_128x32x128_bs256 0.44 4.29 4.28 4.16
ImageMLPAE_Mnist_128x32x128_bs128 0.19 2.54 2.70 2.22
RNNLM_lm1b32k_Patch32_LSTM256_Embed128 23.96 53.42 53.38 52.86
RNNLM_wikipediaen32k_Patch32_LSTM256_Embed128 24.29 53.47 53.81 53.38
TransformerLM_LM1B_MultiRuntime_0 0.71 8.30 8.71 7.10
TransformerLM_LM1B_MultiRuntime_2 0.99 14.50 15.46 12.94
TransformerLM_LM1B_MultiRuntime_5 1.26 10.93 11.42 9.88
LOpt_AdafacMLPLOpt_FashionMnist_50 65.12 67.31 68.16 67.06
Table 7: Runtime per step for each task (ms). Celo is faster in terms of wallclock time per step than previous state-of-the-art learned optimizers such as VeLO (Metz et al., 2022b) as well as VeLO with smaller hidden size (VeLO-S from §5). However, compared to Adam, Celo has additional overhead depending on the evaluation task. Notably, as tasks become more computationally expensive, this overhead diminishes, as discussed in §A.5.
A.6 MLCommons AlgoPerf evaluation
We additionally evaluate on a few standard ML Commons tasks from AlgoPerf benchmark (Dahl et al., 2023) which are much larger and completely out-of-distribution for meta-trained optimizers in this work. Figures 10 and 10 show training and validation curves respectively, comparing Celo-Adam (Celo scheduler with Adam update rule discussed in Table 3) with Schedule-Free AdamW, the winning entry in the self-tuning track of AlgoPerf 2024 competition as well as tuned NAdamW baseline provided by the AlgoPerf team (Kasimbeg et al., 2025). For baselines, we directly use the results data from the AlgoPerf 2024 competition. AlgoPerf is a time-to-result benchmark, where runs are halted either upon reaching the task’s validation target or upon timing out. As evident from the results, Celo-Adam manages to optimize these relatively large-scale unseen workloads despite being meta-trained on small image MLP tasks (§A.7). However, the final validation performance of Celo-Adam still lags behind the tuned baselines and it fails to hit validation targets earlier as evident from the validation curves (Figure 10). Morever, we find that Celo, pre-trained on simple image MLP tasks considered in this work, is unstable for MLCommons tasks, hence we omit it from Figures 10 and 10. This highlights that learning a robust update rule is critical to scale fully learned optimizers to large-scale workloads since learned schedulers can generalize, as evident from Celo-Adam results, even from limited meta-training. We defer the scaling up of meta-training for Celo to future work, which we believe is the right direction to make it ready for standard use-cases especially given that our proposed recipe is compute-efficient and scalable (A.4).
A.7 Meta-training tasks
@@ A.8 Evaluation tasks
Figure 9: MLCommons AlgoPerf train curves.
Figure 10: MLCommons AlgoPerf validation curves. We compare Celo-Adam with winner in the self-tuning track of AlgoPerf 2024 competition Schedule-Free AdamW and prize qualification baseline NAdamW which is heavily tuned (Kasimbeg et al., 2025; Dahl et al., 2023). Although Celo-Adam is just meta-trained on small image MLP tasks (§A.7), it manages to optimize much larger tasks, in some cases, even better than Schedule-Free AdamW (ImageNet ResNet tasks). However, it still lags behind baselines in terms of final performance, a gap that can be potentially covered by improving and scaling up meta-training.
@@ A.9 Evaluation plots
A.10 Evaluation plots with task augmentation